JOHNSON & JOHNSON AND SUBSIDIARIES CONSOLIDATED STATEMENTS OF EARNINGS (Dollars and Shares in Millions Except Per Share AmounJOHNSON & JOHNSON AND SUBSIDIARIES CONSOLIDATED STATEMENTS OF EQUITY (Dollars in Millions) (Note 1) Accumulated Other Compreh(Dollars In millions) 2018 51 Cost of products sold Selling, marketing and administrative expenses Research and development e1. Summary of Significant Accounting Policies Principles of Consolidation The consolidated financial statements include the apresented in the financial statements. The Companys operating leases will result in the recognition of additional assets andcustomers for retumed goods. The Companys sales retums reserves are accounted for in accordance with the U.S. GAAP guidance2,200 Goodwill as of December 30, 2018 and December 31, 2017, as allocated by segment of business, was as follows: (Dollars iOther Cost of BAD (Income) The following table is a summary of the activity related to derivatives and hedges for the fiscalQuestion 11 1.25 points Save Ansv Calculate Johnson & Johnsons return on assets ratio (net earnings / total assets) for theQuestion 12 Johnson & Johnson increased which account when it originally issued shares of its stock for cash? a. Retained earQuestion 13 At the most recent balance sheet date, Johnson & Johnsons shareholder claims to the companys assets totaled (inQuestion 14 At the most recent balance sheet date, Johnson & Johnsons contributed capital totaled (in millions): a. $3,120 0Question 15 Johnson & Johnsons undistributed earnings at the most recent balance sheet date was (in millions): a $94.114 b.
Answer #1

Q11
Return on Assets = Net Income / Total Assets
2018 = 15297 / 152954 = 10%
2017 = 1300 / 157303 = 0.83%

Since ROA increased over the period,

Answer is a. used its assets more efficiently to generate earnings

Q12.
Answer is c. Common Stock
For issue of stock for cash, common stock is credited

Q13.
Answer is a. $59752 millions

Shareholders Claim to assets is Shareholders Equity

Q14.
Answer is a. $3120

It is value of common stock

Q15.
Answer is b. $106216
It is equal to retained earnings